OpenStudy (anonymous):

which ordered pair is the solution to the system below x+3y=7 or x +2y=10

OpenStudy (anonymous):

x=2n y=5-n

OpenStudy (saifoo.khan):

x+3y =7 x +2y=10 --------- Multiply the second equation by -1 x + 3y = 7 -x-2y = -10 ======== 0 + y = -3 y = -3 Now substitute this value in any equation to get the value for x.

OpenStudy (anonymous):

what

OpenStudy (saifoo.khan):

?